# Sparse Auto-Encoder Interprets Linguistic Features in Large Language Models

### v1

@@ 4.1 Experiment Setup
For feature analysis, we select $18$ canonical linguistic phenomena from the renowned textbook An Introduction to Language Fromkin et al. (2017) as experimental features. The selected linguistic features include: Phonetics/Phonology: vowels, fricatives, and stress; Morphology: noun pluralization and past tense formation (inflectional), and agentive suffixes (derivational); Syntax: verb valency (transitive/intransitive), inversion, subject-copula-predicate constructions, and object-complement structures; Semantics: simile, metaphor, contrast, sequence, and causality; Pragmatics: politeness expressions.
We construct sentences that significantly contain $18$ selected linguistic features, called SAELing-Data. Additionally, we create minimal pairs and counterfactual datasets for these sentences, called SAELing-Data-MN and SAELing-Data-CF.
@@ 4.2 Main Results
The main experiments to verify that SAELing finds linguistic features in the SAE space and intervening on these features is effective.
4.2.1 Feature Extraction and Analysis
Feature ID M-Pairs C-Fact FRC
PS PN PS PN
Phonetics
Sibilant 7L230243 $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Vowel 7L82620 $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Phonology
Stress 9L156767 $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
@@ 4.2.1 Feature Extraction and Analysis
Past-Tense 8L4016 $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
Agentive Suffix 12L199760 $100.0$ $95.0$ $100.0$ $92.5$ $96.9$
Plural Noun 12L160227 $100.0$ $97.5$ $100.0$ $100.0$ $99.4$
@@
Intransitive Verb 17L63597 / / $92.5$ $97.5$ $94.9$
Transitive Verb 17L174515 / / $100.0$ $90.0$ $94.7$
Linking Verb 18L61112 / / $90.0$ $95.0$ $92.4$
Inverse 21L17802 / / $100.0$ $100.0$ $100.0$
Genitive 20L259762 $100.0$ $100.0$ $100.0$ $100.0$ $100.0$
@@
Discourse Marker 27L173147 $92.5$ $92.5$ $92.5$ $92.5$ $92.5$
Politeness 31L578 / / $100.0$ $96.3$ $98.1$
Table 1: Feature analysis results. Probability of Necessity (PN) and Probability of Sufficiency (PS) of the extracted linguistic features (Feature, layer, ID) on Minimal-Pair and Counter-Factual datasets, and Feature Representation Confidence (FRC) for each feature.
We input the sentences from SAELing-Data into Llama-3.1-8B and pass the neuron activation distributions after batch normalization through the corresponding SAE layers. We then encode the activation distributions of base vectors at each token in each sentence. The ratio of sentences activated by each base vector is calculated. These base vectors are ranked by this ratio, and their activation is tested on SAELing-Data-MN and SAELing-Data-CF. Base vectors that are significantly activated in SAELing-Data but almost inactive in Datasets SAELing-Data-MN and SAELing-Data-CF are selected as potential base vectors for each linguistic feature. The detailed process mode of feature extraction can be found in Appendix B.1.
For each linguistic feature’s potential base vectors, we compute their activation in SAELing-Data, SAELing-Data-MN, and SAELing-Data-CF. Using the method described earlier, we calculate the sufficiency and necessity probabilities for each base vector in the minimal pairs and counterfactual datasets, and ultimately compute the FRC score for each feature.
Table 1 shows the activation indices of a representative base vector for each linguistic feature. Due to the varying significance of linguistic features across different sentences, activations corresponding to the base vectors may be lost during the sparse forward propagation process of the SAE, particularly in the TopK activation function mode. We consider a base vector to be strongly related to a linguistic feature if its sufficiency probability exceeds $0.9$. For necessity probability, as existing SAEs still exhibit slightly poly-semanticity issues Olah (2024), there are still a small amount of low activation for base vectors in positions that do not correspond to the linguistic features. We assume these low activations do not affect the quality of the base vector’s representation of the linguistic feature, as we discussed in Appendix B.4.
Overall, the base vectors extracted for features across various linguistic levels show strong correlations. The high quality of phonetic and phonological features indicates that the model contains accurate IPA-related knowledge. The performance of the single feature for metaphor is suboptimal, suggesting that the representation and processing of metaphors may involve more complex mechanisms. Furthermore, FRC values for features in other dimensions mainly exceed $0.9$, demonstrating that the selected typical linguistic features consistently yield highly correlated base vectors.
@@ 4.2.2 Feature Intervention
We select 5 representative ones for intervention experiments. The intervention method involves modifying the activation values of specific base vectors (by index) on a designated SAE layer during forward propagation. We perform two types of intervention: feature enhancement and ablation. With identical input tokens, we set the activation value to $10$ for enhancement and 00 for ablation. We then compare the generated outputs with those from the unmodified SAE model, focusing on the prominence of the target linguistic features.
We find that intervening on a single linguistic feature in one layer does not produce effects that are easily distinguishable by human evaluators. Therefore, we use an LLM (GPT-4o) as a judge Zheng et al. (2023) to assess the prominence of these features in the outputs. For each feature, we conduct $50$ experiments and calculate the probabilities of successful enhancement and ablation, i.e., increased and decreased feature prominence, respectively.
In addition, we randomly select 50 base vector indices from the intervention layer and conduct enhancement and ablation experiments under the same conditions as a control. The success rates in the control group are not around $0.5$; typically, the enhancement success rate is below $0.5$ while the ablation success rate is above $0.5$. This discrepancy may stem from the intervention affecting the model’s output quality, thereby interfering with the proxy LLM’s judgment.
We compute the efficacy of the selected base vectors in both experiments and calculate the FIC value and show the results in Table 2. Our results show that enhancement experiments yield significantly better effects than ablation experiments, with all features demonstrating marked enhancement effects. In the ablation experiments, the politeness feature shows relatively good performance, while other features are less affected; the simile feature does not yield the desired ablation effect. This may be because multiple features in the model control the same linguistic phenomenon. Enhancement interventions have a larger impact on the model, whereas ablation of a single feature may be compensated by other features, leading to suboptimal ablation outcomes. Overall, all 5 features exhibit clear causality in the intervention experiments.
@@ 4.3 Analysis
We further conduct analytical experiments to explore the property of SAELing.
@@ 4.3.1 Combined Intervention
We select two linguistic features—simile and politeness—from layer 26. Each feature has four highly related base vectors in this layer. We increase the number of intervened features from one to four. In each experiment, we randomly chose the specified number of base vectors from the four. We used GPT-4o to assess the prominence of the targeted linguistic feature in the generated outputs. For each feature, we conducted $200$ enhancement experiments and $200$ ablation experiments. We also perform control experiments in layer 26 by randomly selecting a set number of base vectors to intervene.
@@
4.3.2 Deep Semantics Processing
Deep semantics refers to the underlying meaning structures that extend beyond surface-level syntax and lexical definitions. It captures implicit relationships and conceptual associations within language. We conduct experiments to show that SAE
Reference and metaphor exemplify deep semantics by utilizing cognitive mappings and contextual dependencies to convey meaning beyond explicit expression. We conduct experiments on reference and metaphor at the sixth and fifteenth layers respectively. From the results shown on Figure 4, we observe the following:
Metaphor. In the metaphor sentence, our experimental statements contain only the vehicle (fire) and not the tenor (sun). In the 6th layer, the base vector corresponding to the vehicle is activated, while the base vector for the tenor remains inactive. In the 15th layer, the activation of the vehicle’s base vector decreases, while the base vector for the tenor becomes activated. This suggests that as the model moves to deeper layers, the vehicle maps to the target domain and generates the deep semantics of the tenor, even without the tenor in the context.
4.3.3 Cross-layer Activations
Figure 5: Activation value distributions of deep semantic corresponding features in the 6th and 15th layers for anaphoric and metaphor example sentences.
In different layers of the model, we identify distinct base vectors that are activated by the same linguistic feature. To validate this phenomenon, we select a standard base vector from a given layer for each linguistic feature and apply the corresponding SAE to other layers of Llama-3.1-8B. Encoding the hidden states of these layers with the cross-layer SAE, using the same dataset, reveals that in most layers the base vector with the same index as the standard one shows a similar activation pattern on the tokens.
This approach provides an effective tool for observing the cross-layer activation of linguistic features. Figure 5 displays 8 linguistic features. For each feature, the average maximum activation curve and the reciprocal rank curve (used to assess the relative importance of the base vector on the dataset) of the standard base vector are shown for different layers. The distributions indicate that: (1) Linguistic features are widely activated across layers, with only some features failing to activate in the early or late layers; (2) Features that do not activate in the early layers are mostly semantic, while those that do not activate in the later layers are mainly syntactic, possibly reflecting the model’s functional division across layers; (3) Activation curves typically follow a “rise–plateau–fall” pattern, whereas the reciprocal rank curves often exhibit several sharp peaks. This suggests that each linguistic operation has its own active processing range within the model, although its relative importance may vary across layers.
Figure 6: Distribution of average maximum activation values and reciprocal ranks (i.e., the reciprocal of the base vector ranking) for seven typical linguistic features across layers in the dataset.
We further examine the cross-layer activation pattern at the sentence level using the metaphor feature as an example in Figure 6. We select the standard metaphor feature from layer 26 and analyze its activation distribution in layers 2, 5, 10, 14, 19, 24, and 29. In both example sentences, layers 2 and 29 show no token activations, suggesting that the model’s initial and final layers may not process metaphors. Layers 5, 10, and 14 appear to be involved in preliminary metaphor processing: in the first sentence, the feature intermittently activated the token following the cue word “like”; in the second sentence, the feature first activated a fixed structure (“as as”), then activated the second “as” and the subsequent token. Around layer 19, the activation extends to the cue word and the entire vehicle, while around layer 24, the activation recedes back to the cue word. These results indicate more complex internal mechanisms for metaphor processing and demonstrate that SAE-based feature extraction is a valuable tool for further exploring the model’s internal linguistic mechanisms.
@@ 4.3.4 Case Study for Intervention
We conduct manual case study on the generated contents after intervening on one identified simile-related base vector. We show cases in Table 3.
In Case #1, the prompt is “Generate a sentence describing winter”, which does not explicitly include the target linguistic feature. We find that after enhancing the simile-related base vector, the LLM turns to use simile. We can also find that the descriptive and imagistic quality of the default output is stronger than in the ablation results, which indicate that the simile-related base vector is also responsible for vividness.
Case #2 uses the prompt “Generate a sentence using a simile to describe love”, with explicit requirement for using simile to generate the sentence. When the simile-related base vector is ablated, the LLMs turn to use straightforward descriptions without using similes. Meanwhile, when enhancing the simile-related base vector, the LLMs continue to generate sentences with simile. We show more intervention cases in Appendix C.1.
@@ 5 Conclusion
This work addresses two key challenges in interpreting linguistic mechanisms in LLMs: coarse interpretation granularity and insufficient causal analysis. We introduce SAELing, which uses sparse autoencoders (SAE) for fine-grained feature extraction, overcoming poly-semanticity in traditional methods. SAELing also verifies causal relationships by intervening on features, enabling more precise control over model behavior. Our approach reveals that LLMs encode structured linguistic knowledge and offers a robust framework for steering their outputs.
@@ 6 Limitations
Our work has several limitations in terms of dataset size, number of features, intervention effects, and cross-layer analysis. Regarding datasets, each feature in our study is constructed with approximately 160 sentences. In the future, the dataset can be further expanded to serve as a benchmark for evaluating the language mechanism interpretability of the SAE framework. Concerning the number of selected features, we choose 18 representative linguistic features from various theoretical linguistic dimensions. This selection sufficiently demonstrates the effectiveness of our method across different linguistic levels; however, to construct a complete and comprehensive language mechanism system, our approach can be extended to extract a larger number of linguistic features. Achieving this extension will require further work or the development of automated feature extraction methods. In terms of intervention effects, our experiments show statistically significant effects for linguistic feature interventions, yet the effect and stability of each case are still inferior to fine-tuning methods. This issue calls for further research and refinement of SAE-based intervention methods. Finally, regarding cross-layer analysis, our experiments illustrate the cross-layer mechanism of linguistic features, revealing the potential of our method to explain how large language models process and transmit linguistic information across layers. However, we do not conduct large-scale experiments and inductive analyses in this area, which represents an extension of our method that remains to be explored in future work.
@@ 7 Ethical Considerations
SAELing is intended to be used as a method for analyzing the mechanisms of large language models.

### v2

@@ 4.1 Experiment Setup
For linguistic feature analysis, we select a total of 145 linguistic features—99 in English and 46 in Chinese—spanning four core domains: morphology, syntax, semantics, and pragmatics. For each feature, we generate 50 sentences that exhibit the feature and 50 corresponding counterfactual sentences, yielding a large-scale dataset for systematic feature extraction and analysis.
@@ 4.2 Main Results
The main experiments to verify that LinguaLens finds systematic linguistic features in SAE space and intervening on these features is effective.
4.2.1 Feature Extraction
Lang PS PN FRC Actm
0 8 15 24 30
@@ 4.2.1 Feature Extraction
CH $0.61$ $0.70$ $0.64$ $0.01$ $0.19$ $0.29$ $0.52$ $1.36$
EN $0.73$ $0.80$ $0.75$ $0.03$ $0.35$ $0.49$ $1.02$ $1.89$
@@
CH $0.84$ $0.90$ $0.86$ $0.20$ $0.50$ $0.95$ $2.32$ $3.37$
EN $0.79$ $0.87$ $0.82$ $0.12$ $0.35$ $0.68$ $1.66$ $2.59$
@@
CH $0.69$ $0.74$ $0.70$ $0.06$ $0.25$ $0.42$ $1.03$ $1.56$
EN $0.77$ $0.83$ $0.79$ $0.13$ $0.27$ $0.52$ $1.33$ $2.03$
Table 1: Extracted feature analysis. The mean representation metrics (PS, PN, FRC, and max activation) for morphological, syntactic, semantic, and pragmatic features in both Chinese and English.
We feed the sentences from LinguaLens-Data into Llama-3.1-8B and, after batch normalization, pass the resulting neuron activation distributions through the corresponding SAE layers. For each sentence and each token, we then encode its activation distribution over the SAE base vectors at every layer. As described in the Methods, we compute the probability of sufficiency (PS), probability of necessity (PN), and FRC for each base vector on the counterfactual datasets at each layer, rank the base vectors by FRC, and use GPT-4o to select the feature-corresponding vectors based on their activation patterns. For a detailed description of the feature-extraction procedure, see Appendix B.
To evaluate how well a given layer represents a particular linguistic feature, we calculate the arithmetic mean of PS, PN, and FRC for the selected base vectors, as well as their average maximum activation on the positive examples (if more than three vectors are identified, we select the top three by FRC).
Table 1 reports, for layers 0, 8, 15, and 30, the mean representation metrics (PS, PN, FRC, and max activation) for morphological, syntactic, semantic, and pragmatic features in both Chinese and English.
Overall, at these representative layers, the base vectors extracted for features across different linguistic levels exhibit strong correlations. From layer 0 to layer 30, the average maximum activation exhibits a monotonic increase. Across the four linguistic domains, syntactic features attain the highest mean maximum activations, followed by semantic and pragmatic features, while morphological features remain lowest. Moreover, substantial discrepancies emerge between the average maximum activations for Chinese and English features, indicating potential differences in the model’s internal representations and processing mechanisms for the two languages. These cross-lingual variations will be explored in greater depth in subsequent analyses.
@@ 4.2.2 Feature Intervention
We select 6 representative features for the intervention experiments. The intervention method involves modifying the activation values of specific base vectors (by index) within a designated SAE layer during forward propagation. We perform two types of intervention: feature enhancement and ablation. Under identical input token conditions, we set the activation value to 10 for enhancement and to 0 for ablation. We then compare the outputs generated after intervention with those from the unmodified SAE model, focusing on the prominence of the target linguistic features.
We find that intervening on a single linguistic base vector in one layer does not produce effects easily distinguishable by human evaluators. Therefore, we employ an LLM (GPT-4o) as a judge Zheng et al. (2023) to assess feature prominence in the outputs. For each feature, we conduct 50 experiments and calculate the enhancement success rate and ablation success rate—that is, the probabilities of increased and decreased feature prominence, respectively. Furthermore, for each linguistic feature, we select three base vectors with the highest FRC as representatives for intervention and compute the average results across these three interventions.
In addition, we randomly select 50 base vector indices from the same layer and perform enhancement and ablation experiments under the same conditions as a control. The control group’s success rates do not converge around 0.5; typically, enhancement rates fall below 0.5 while ablation rates exceed 0.5. This discrepancy may arise because the intervention affects overall output quality, thereby confounding the proxy LLM’s judgments.
We compute the efficacy of the selected base vectors in both experiments and derive the FIC values; the results are presented in Table 2.
Our results show that enhancement experiments yield significantly stronger effects than ablation experiments, with all features demonstrating marked enhancement. In ablation experiments, the politeness feature shows relatively good performance, whereas other features are less affected; the simile feature fails to achieve the desired ablation effect. This may be because multiple base vectors collaboratively control the same linguistic phenomenon. Enhancement interventions have a larger impact on the model, while ablating a single feature can be compensated by other vectors, leading to suboptimal ablation outcomes. Overall, all 6 features exhibit clear causal effects in the intervention experiments.
@@ 4.3 Analysis
We further conduct analytical experiments to explore the properties of LinguaLens.
4.3.1 Multilingual Analysis
Figure 3: Heatmap of the overlap between Chinese and English feature sets across the SAE basis vectors at each of 32 layers. The horizontal axis groups Chinese and English features with analogous form and function—ordered by morphology, syntax, semantics, and pragmatics—while the vertical axis indexes the model layers. Darker red indicates greater overlap.
We investigate the multilingual mechanisms of the model. We select Chinese and English as test languages and choose $24$ sets of feature collections representing the same linguistic functions, including set $2$ of morphological features, set $11$ of syntactic features, set $6$ of semantic features, and set $5$ of pragmatic features. We test the degree of overlap between the latent-space basis vectors activated internally by the model when representing these features in Chinese vs. English. The overlap for layer $i$ is computed as follows: let the set of English basis vectors for the feature at layer $i$ be $\mathrm{Eng}_{i}$, and the corresponding Chinese set be $\mathrm{Chi}_{i}$, then
$\mathrm{overlap}_{i}=\frac{|\mathrm{Eng}_{i}\cap\mathrm{Chi}_{i}|}{|\mathrm{Eng}_{i}|}.$
After computing the overlap for each layer, we aggregate the overlap rates for all feature pairs across layers into a matrix and visualize it with a heatmap. The results yield the following conclusions:
Linguistic Levels.
The overlap between Chinese and English features is greater at the semantic and pragmatic levels, but lower at the morphological and syntactic levels, indicating that cross-lingual linguistic knowledge representations are primarily manifested at the semantic and pragmatic levels.
Model Layers.
The overlap is higher in the first 16 layers and lower in the latter 16 layers, suggesting that the deep semantic computations in the model’s upper layers are less correlated with cross-lingual universal linguistic features.
LinguaLens demonstrates its potential for analyzing models’ cross-lingual knowledge representations, laying the foundation for further analysis and transfer in low-resource languages.
4.3.2 Deep Semantics Processing
Deep semantics refers to the underlying meaning structures that extend beyond surface-level syntax and lexical definitions. It captures implicit relationships and conceptual associations within language. We conduct experiments to show that SAE can interpret the mechanism of deep semantics.
Reference and metaphor exemplify deep semantics by utilizing cognitive mappings and contextual dependencies to convey meaning beyond explicit expression. We conduct experiments on reference and metaphor at the sixth and fifteenth layers respectively. From the results shown in Figure 4, we observe the following:
Metaphor. In the metaphor sentence, only the vehicle (fire) is included, while the tenor (sun) is omitted. In the 6th layer, the base vector corresponding to the vehicle is activated, while the base vector for the tenor remains inactive. In the 15th layer, the activation of the vehicle’s base vector decreases, while the base vector for the tenor becomes activated. This suggests that as the model moves to deeper layers, the vehicle maps to the target domain and generates the deep semantics of the tenor, even without the tenor in the context.
4.3.3 Cross-layer Functions
S. L. Descrip. Top 10 Features
I 0–2 Mor.&BS past tense, verbal suffix, adjectival suffix, noun plural, possessive genitive, linking verb, passive voice, anaphor, extraposition, factives
II 3–8 CS&EP elliptical sentences, relative clauses, subject auxiliary inversion, emphatic structure, existential quantifiers, coordination, cleft sentences, light verbs, reduplication, metaphor
III 9–16 Di.&Prag. interrogative, tag questions, subjunctive mood, optative, turn taking, discourse markers, intensifiers, euphemism, politeness, coordination
IV 17–31 DS&RS personification, synecdoche, metaphor, expressive pragmatics, imperative, directive pragmatics, topic comment, representative pragmatics, euphemism, politeness
Table 3: The four hierarchical stages of the model’s linguistic functions. For each stage, the ten features with the highest activation frequency and largest activation values are displayed. S., L. and Descrip. stand for Stages, Layers and Descriptions , respectively.
We further investigate how the model’s linguistic functions distribute across layers. We assemble 50 English sentences—drawn both from classic texts and manually crafted—to cover a broad range of linguistic phenomena. For each sentence, we record every activated basis vector and its activation value at all 32 layers. By comparing these activated vectors against our pre-compiled dictionary of linguistic feature vectors and computing their overlap, we determine which linguistic functions each layer encodes. We then identify, for every layer, the 10 features with the highest activation frequency and magnitude. Aggregating results over all 50 sentences, we distill four processing stages as Table 3:
Stage I (layers 0–2) primarily encodes morphology and basic syntax features (abbreviated as Mor.&BS). Stage II (layers 3–8) introduces complex syntactic phenomena and early pragmatic cues (abbreviated as CS&EP). Stage III (layers 9–16) focuses on discourse and pragmatic markers (abbreviated as Di.&Prag.). Stage IV (layers 17–31) integrates deep semantics and rhetorical structure (abbreviated as DS&RS).
These results reveal the functional division of labor across layers: lower layers handle morphology and syntax, middle layers capture pragmatics and context, and upper layers perform holistic semantic computation.
@@ 4.3.4 Combined Intervention
We select two linguistic features—simile and politeness—from layer 26. Each feature has four highly related base vectors in this layer. We increase the number of intervened base vectors from one to four. In each experiment, we randomly chose the specified number of base vectors from the four. We use GPT-4o to assess the prominence of the targeted linguistic feature in the generated outputs. For each feature, we conduct $200$ enhancement experiments and $200$ ablation experiments. We also perform control experiments by randomly selecting a set number of base vectors to intervene.
@@ 4.3.5 Case Study for Intervention
We conduct a manual case study on the generated content after intervening on one identified simile-related base vector. We present cases in Table 4.
In Case #1, the prompt is “Generate a sentence describing winter”, which does not explicitly include the target linguistic feature. We find that after enhancing the simile-related base vector, the LLM turns to using a simile. We can also find that the descriptive and imagistic quality of the default output is stronger than in the ablation results, which indicates that the simile-related base vector is also responsible for vividness.
Case #2 uses the prompt “Generate a sentence using a simile to describe love”, with explicit requirement for using a simile to generate the sentence. When the simile-related base vector is ablated, the LLMs turn to use straightforward descriptions without using similes. Meanwhile, when enhancing the simile-related base vector, the LLMs continue to generate sentences with similes. We show more intervention cases in Appendix D.1.
@@ 5 Conclusion
We propose LinguaLens, a method to help solute the coarse-granularity problem in linguistic mechanistic studies and a means to enable large-scale, systematic study of linguistic mechanisms in LLMs. Our approach comprises two key components: (1) a comprehensive counterfactual dataset of linguistic features, and (2) an SAE–based framework for feature extraction, together with causal validation through interventions. Using LinguaLens, we conduct an in-depth analysis of the model’s multilingual representation mechanisms and the cross-layer distribution of linguistic functions. Our results demonstrate that LLMs inherently encode structured linguistic knowledge and provide a robust framework for steering model outputs.
@@ 6 Limitations
Our work has several limitations in terms of dataset size, feature count, experimental model, and intervention effects.
In datasets, each linguistic feature is constructed from approximately 50 pairs of example and counterfactual sentences. In the future, this dataset can be further expanded to serve as a standard benchmark for linguistic-mechanism interpretability.
In feature count, we select 145 representative linguistic features from various theoretical dimensions to validate our method at scale across different layers; however, building a fully comprehensive linguistic-mechanism system requires extending to even more features, which will depend on further work.
In experimental model, due to computational constraints we use Llama-3.1-8B for all experiments. In future work, our dataset and analytical framework can be applied to a wider variety of architectures and larger models for deeper linguistic-mechanism analysis.
In intervention effects, although our experiments show statistically significant effects from feature-based interventions, the efficacy and stability of single interventions remain inferior to conventional fine-tuning techniques. This shortcoming calls for further research to refine SAE-based intervention methods.
@@ 7 Ethical Considerations
LinguaLens is intended to be used as a method for analyzing the mechanisms of large language models.
@@
Acknowledgment
This work is supported by Beijing Natural Science Foundation (L243006), National Natural Science Foundation of China (62476150) and the Tsinghua University Disruptive Innovation Talent Development Program. We would also like to sincerely thank Prof. Weidong Zhan from Peking University, who discussed with us and provided many valuable insights in linguistics.